OpenStudy (anonymous):

Solve the inequality: 3(2+t)_> 15-2t Please show steps

OpenStudy (tkhunny):

Didn't we already do this? Use the distributive property to remove the parentheses. 3(2+t) = 6 + 3t

OpenStudy (anonymous):

it is not complete which i need. I also need steps shown. :/

OpenStudy (anonymous):

\[6+3t \ge25-2t\] \[3t+2t \ge25-6\]

OpenStudy (anonymous):

\[5t \ge19\]

OpenStudy (anonymous):

\[t \ge19/5\]

OpenStudy (tkhunny):

It is YOUR homework. You need to DO the steps, not just SEE them. SHOW your work! juniro won't be there on your exam.

OpenStudy (tkhunny):

It's also kind of misleading, since juniro very nicely worked the wrong problem. Now, it is your turn to work the right problem. Let's see what you get and HOW you get it. Show us the steps!

OpenStudy (anonymous):

I am stuck on one problem, sorry for asking for some help. I understand MY homework I just dont understand this one, and I will do fine on my "exam" even though we will not be having them.

OpenStudy (tkhunny):

Prove it! Are you going to argue with me or show your work? What's the impediment on this one? Never done an inequality? What is it that seems more difficult than other similar problems that you have solved correctly?

OpenStudy (tkhunny):

Can you solve this? 3(2+t) = 15-2t

OpenStudy (anonymous):

I am not going to sit here and argue when I can find help elsewhere. I'm tired and I need help not to "prove" I can solve an inequality.

OpenStudy (tkhunny):

You will fail your exam with that attitude. Good luck. I wish I had better news, but you won't let me. Come back when you're ready.

OpenStudy (tkhunny):

Having someone else work the wrong problem for you is not very instructive. Seriously, though, good luck with your vicarious studies.

OpenStudy (anonymous):

As I have said before we do not have exams and I have never failed any test or homework assignment. Only once on an English paper because my friend copied my essay. I do not need to be "ready" to come onto a homework help website lol

OpenStudy (tkhunny):

You do need to be ready to SHOW YOUR WORK. If you have nothing to offer, you are in the wrong place. As you said, there are plenty of places that will do the work for you. You call this "help". I disagree. Good Day.

OpenStudy (anonymous):

lol goodnight sorry I got under your skin